A criminal is escaping across a rooftop and runs off the roof horizontally at a speed of 5.2 m/s, hoping to land on the roof of an adjacent building. Air resistance is negligible. The horizontal distance between the two buildings is D, and the roof of the adjacent building is 2.0 m below the jumping-off point. Find the maximum value for D.
1
Expert's answer
2013-09-27T08:49:19-0400
what time is needed to pass 2 m verticaly?


s=gt^2/2, t = sqrt(2s/g)=0.64s


in 0.64 seconds he can pass horizontally


D=t*v = 0.64*5.2 = 3.32m


So the maximum for D is 3.32 m